by TEI TEI

12/31/09 – There are minor control issues with the economizer in question but the building and controls are working as designed. TTC archived on 12/31/09.

10/20/09 – Kevin has JCI coming out to change return air humidity to zone average humidity on RTU-1 on Thursday. TTC may attend. Three items remain on the service punchlist. TTC contacted Brett to get answers about economizer control,

09/22/09 – I believe the final controls punchlist items to be complete is to acquire the software CD’s and backups of the current settings.

08/25/09 – Building is overcool in mornings. JCI is supposed to come out and fix timing and response loops. Many emails and phone calls have been made trying to get this done.

07/29/09 – Today we received updated As-Builts on the controls for the building from Chuck Vedder. I forwarded them to Kevin Maue for his keeping. I am also printing out a full set of 11×17’s to keep in the basement of the MRDD building again for our use. Everything seems to be running really really well.

06/23/09 – Triton, JCI and Carrier have finally been to the building enough to help us change the programming on the unit to get it to coordinate with the JCI system. This project may finally be finished. Things are working well in both buildings.

02/02/09 – Triton is reporting that Chuck Vetter will finalize this project for 1520 Madison. TTC contacted TJ Williams for the smoke detector work on the 1956 unit. TTC sent a summary email for controls to Brett at Triton for final completion items. The unit is finally running well but the building pressurization & the duct heaters are not working quite right yet.

10/16/08 – Grady says Triton will put together a final sequence that incorporate everything that is really working on the project and the way it works.

09/26/08 – Const Mtg 2 is over. TEI responded to TJ Williams RFi about another light in the hallway. TEI is approving the PHB Honeywell controller for the HVAC unit.

08/25/08 – Triton emailed today to discuss remaining closeout items. TTC explained about OA sensor.

08/22/08 – Triton is ordering a new board for Boiler 2. Triton relieved the pressure on the hot water system to prevent relief valve issues. Triton may have to get TJ to troubleshoot the overloading problem on the recirculation pumps. Kevin emailed that the OA sensor was flaky again.

07/29/08 – Brett Harpel of Triton discussed final control sequences with JCI. Modifications to system were made a month ago and again last week. Building is running fine. KM wants TEI to bless the control system as it meets design intent of the specs. JCI replied stating TEI sequences were in fact in place on RTU-2 but local overrides were keeping the system from operating as designed on that day.

Triton asked to switch the disconnects serving the RCP to troubleshoot low voltages problems at Boiler-2.

Triton to test expansion tank pressure to see if it is causing over pressurization.